The systolic blood pressure of adults in the USA is nearly normally distributed with a mean of 122 and standard deviation of 22 .
Someone qualifies as having Stage 2 high blood pressure if their systolic blood pressure is 160 or higher.
a. Around what percentage of adults in the USA have stage 2 high blood pressure? Give your answer rounded to two decimal places.
b. If you sampled 2000 people, how many would you expect to have BP> 160? Give your answer to the nearest person. Note: I had a bit of an issue encoding rounded answers, so try rounding both up and down if there's an issue!
c. Stage 1 high BP is specified as systolic BP between 140 and 160. What percentage of adults in the US qualify for stage 1?
d. Your doctor tells you you are in the 30th percentile for blood pressure among US adults. What is your systolic BP? Round to 2 decimal places.

Answer by Boreal(15235)

z=(x-mean)/sd
=(160-122)/22=38/22=1.73
probability of z>1.73 is 0.0418 or 4.18%
for 2000, using the rounded answer the result is 83.63 or 84 people
Not rounding even the z-value, the answer is 4.21% but this is still 84 people.

For 140 to 160, the first is (140-122)/22 or 18/22 or z of +0.82
For 160, z is +1.73 as above
the probability of z between those two values is 0.1643 or 16.4%
30th percentile is z of -0.52
-0.52=(x-mean)/22
-11.44=x-122
x=110.56 mm Hg